# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v2

@@ A.2 Proof of global statistical Error Bound
$\begin{split}&\quad-\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}+\frac{\lambda}{2}\sum_{i}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}\\ &=\sum_{i}\left(\frac{\lambda}{2}p_{i}\|\tilde{\bm{\delta}}^{(i)}\|^{2}-\rho\frac{p_{i}}{\sqrt{n_{i}}}\|\tilde{\bm{\delta}}^{(i)}\|_{2}\right)\\ &\geq-\frac{\rho^{2}}{2\lambda}\sum_{i}{\frac{p_{i}}{n_{i}}},\end{split}$ (40)
where we choose the optimal $\|\tilde{\bm{\delta}}^{(i)}\|_{2}^{\star}$ as $\frac{\rho}{\lambda\sqrt{n_{i}}}$.
Plugging (LABEL:global:_case2_opt5) into (39), it yields
$0\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\frac{\rho^{2}}{2\lambda}\sum_{i}\frac{p_{i}}{n_{i}}-R\rho\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}-\frac{\lambda}{2}R^{2}.$ (41)
Then we maximize the RHS of (41) w.r.t $\lambda$, the optimal solution of $\lambda$ is $\frac{\rho}{R}\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$. We add the result into (41) and it yields
00 0 $\geq\mu\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}-\rho R\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}-\rho R\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}.$ (42)
We can regard (42) as a function of $\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and induce
$\begin{split}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}&\leq\frac{1}{2\mu}\left\{\rho\sqrt{\sum_{i}\frac{p_{i}^{2}}{n_{i}}}+\left(\rho^{2}\sum_{i}\frac{p_{i}^{2}}{n_{i}}+4\mu\rho R\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)\right)^{\frac{1}{2}}\right\}.\end{split}$ (43)
Thus
$\begin{split}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\frac{\rho^{2}}{\mu^{2}}\sum_{i}\frac{p^{2}_{i}}{n_{i}}+\frac{2\rho R}{\mu}\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right).\end{split}$ (44)
Combine two-case arguments, we have:
@@
$\Rightarrow\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\right)^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}&R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}\\ \frac{\rho^{2}}{\mu^{2}}\sum_{i}\frac{p^{2}_{i}}{n_{i}}+\frac{2\rho R}{\mu}\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)&R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}.\end{cases}$ (45)
@@
$\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq C_{1}\sum_{i}\frac{p_{i}^{2}}{n_{i}}+C_{2}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\wedge R\left(\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}+\sum_{i}\frac{p_{i}}{\sqrt{n_{i}}}\right)\right),$ (46)
@@
$\begin{split}C_{1}&=\frac{\rho^{2}}{\mu^{2}},\\ C_{2}&=\left(\frac{(1+\sqrt{1+2\mu})\rho}{2\mu}\right)^{2}\vee\frac{2\rho}{\mu}.\end{split}$ (47)
@@
$\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq C_{1}\frac{1}{N}+C_{2}\left(\frac{1}{n}\wedge\frac{R}{\sqrt{n}}\right).$ (48)
A.3 Proof of local statistical Error Bound
Proof.
Firstly, we will provide some basic results for the local optimality.
To prove the local statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$ for a single client, which yields
@@ A.3 Proof of local statistical Error Bound
For the first term on the RHS, apply the $\mu$-strongly convex of the loss function $\ell$ (Assumption 2)
$\begin{split}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\end{split}$ (51)
@@
Plugging (51) back into (50), it yields
$\begin{split}0&\geq-\sum_{j}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\mu\sum_{j}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (52)
@@
$\begin{split}\sum_{j}\frac{1}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\rho\frac{1}{\sqrt{n_{i}}}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (55)
@@
$0\geq-\rho\frac{1}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (56)
To utilize the statistical global convergence rate shown in Appendix A.2, we need to argue $R$ in the same way. Recalling the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ if data is balanced (48)
$\Rightarrow\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\begin{cases}\frac{C_{2}}{n}&R>\frac{1}{\sqrt{n}}\\ \frac{C_{1}}{N}+\frac{C_{2}R}{{n}^{\frac{1}{2}}}&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (57)
For simplicity, we denote the upper bound of $\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$
Secondly, we need to discuss the upper bound of $\tilde{\Delta}^{(i)}_{\text{stat}}$ with two cases below.
Case 1: For the first case, we consider the statistical heterogeneity is relative, i.e., $R>\frac{1}{\sqrt{n_{i}}}$.
@@
$\begin{split}-\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}&=-\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &=-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}-2\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\geq-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-R^{2}-2R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\end{split}$ (58)
@@
$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ =&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}-\lambda R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (59)
Put (57) into (59) and reorganize
@@
$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}U_{0}^{2}-\frac{\lambda}{2}R^{2}-\lambda R\;U_{0}\\ \geq&\mu\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-\frac{\lambda}{2}\left(U_{0}+R\right)^{2}.\end{split}$ (60)
@@
In addition, recall that in this case we set $\lambda^{\star}=\frac{\rho^{2}}{R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}$ in the statistical convergence analysis of the global model. If we assume data balanced, i.e. $p_{i}=\frac{1}{m}$ and $n_{i}=n,\forall i\in[m]$, it yields $\lambda^{\star}=\frac{1}{nR^{2}}$

### v3

@@ A.3.2 Proof of Global Statistical Error Bound in Theorem 1
$\begin{split}\sum_{i}p_{i}\left(L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\right)&\leq\sum_{i}p_{i}\left(L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})+\frac{\lambda}{2}\|\ \tilde{\bm{w}}_{\text{GT}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\right)\\ &=\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i}).\end{split}$ (36)
Therefore, we obtain
$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(i)}},S_{i}).$ (37)
In addition, applying the smoothness assumption in Assumption 2 yields
$\left|L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})-L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})^{T}({\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}})\right|$ $\leq\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}.$ (38)
Combining (37) and (38), we have
$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})$ (39)
$\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\sum_{i}p_{i}\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})^{\top}({\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}})-\sum_{i}p_{i}\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}|^{2}.$ (40)
Using Cauchy inequality on the second term of the R.H.S.,
$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|_{2}$
$\quad-\sum_{i}p_{i}\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}.$ (41)
On the other hand, note that the optimality condition of $\tilde{\bm{w}}_{\text{GT}}$ is
$\sum_{i}p_{i}\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})=0,$ (42)
Using the strong-convexity assumption in Assumption 3, we can show that
$\sum_{i}p_{i}\left\{L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})-\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})^{T}({\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}})\right\}$
$=\sum_{i}p_{i}\left\{L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})\right\}$
$\geq\sum_{i}p_{i}\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$
$=\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}.$ (43)
Combining the results (A.3.2) and (A.3.2), it yields
$\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})$ $\geq\sum_{i}p_{i}L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})+\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$
$-\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|_{2}-\sum_{i}p_{i}\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}.$ (44)
Reorganizing these terms yields
$\frac{\mu}{2}\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}\leq\sum_{i}p_{i}\left\{\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|_{2}+\frac{L}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\right\}.$ (45)
Using Lemma 4 to bound the term $\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$, we have
$\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$ $\leq\frac{2}{\mu}\sum_{i}p_{i}\left[\frac{2}{\mu+\lambda}+\frac{2L}{(\mu+\lambda)^{2}}\right]\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}$
$=\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}.$ (46)
Note that by the smoothness Assumption and triangle inequality,
$\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}^{2}$ $\leq 2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\|^{2}+2\|\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\|^{2}$
$\leq 2L^{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}+2\|\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\|^{2}.$ (47)
Take expectation w.r.t $S_{i}$ and use Assumption 1 and 4, then we have
$\mathbb{E}[\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}]$ $\leq 2L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]+2\mathbb{E}\left[\left\|\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\nabla L({{\bm{w}}_{\star}^{(i)}},z_{ij})\right\|^{2}\right]$
$\leq 4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n},$ (48)
where the last term comes from (30).
Plugging (A.3.2) back into (A.3.2), it yields
$\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}]$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\left\{4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right\}.$ (49)
Next, we are ready to establish the global error bound. Using the result (49) and Lemma 5, we obtain
$\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\left\{4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right\}$
$\quad+\frac{8\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{8L^{2}}{\mu^{2}}R^{2}.$ (50)
Let’s define $g(\lambda)=\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)$ and reorganize (A.3.2), which yields
$\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]$ $\leq\frac{g(\lambda)\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right)+\frac{8\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{8L^{2}}{\mu^{2}}R^{2}}{1-4L^{2}g(\lambda)}.$ (51)
If we assume
$g(\lambda)$ $\leq\frac{1}{8L^{2}}\land\left(\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\Big{/}\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\right),$ (52)
we have
$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$ $\leq 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right).$ (53)
Furthermore, we can get a simplified condition for $\lambda$ as
$\lambda\geq\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\Big{/}\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\right)\right\}-\mu$ (54)
where $\kappa=\frac{L}{\mu}\geq 1$ and the condition that $\lambda$ is non-negative holds trivially as $\frac{\kappa L^{2}}{\mu}=\kappa^{2}L\geq\mu$.
Combine two-case arguments, we have
@@
$\Rightarrow\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}&R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}\\ 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)&R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}.\end{cases}$ (55)
@@
$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\wedge R^{2}),$ (56)
@@
$\begin{split}C_{1}&=32\frac{\rho^{2}}{\mu^{2}},\\ C_{2}&=\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\vee\frac{32L^{2}}{\mu^{2}}.\end{split}$ (57)
@@
$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (58)
as claimed.
A.3.3 Proof of Local Statistical Error Bound in Theorem 1
Case 1: first we consider the case when $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
To prove the local statistical error bound, we start with the optimality condition of ${\tilde{\bm{w}}^{(i)}}$ and ${\tilde{\bm{w}}^{(g)}}$ for a single client, which yields that for $i\in[m]$,
@@ A.3.3 Proof of Local Statistical Error Bound in Theorem 1
For the first term on the RHS, applying the $\mu$-strongly convex of the loss function $\ell$ (cf. Assumption 3) yields
$\begin{split}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left(\ell({\tilde{\bm{w}}^{(i)}},z_{ij})-\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\right)&\geq-\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{j\in[n_{i}]}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2},\end{split}$ (61)
@@
Pluggging (61) back into (60), it yields
$\begin{split}0&\geq-\sum_{j}\frac{1}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{j}\frac{1}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.\end{split}$ (62)
@@
$\begin{split}\sum_{j}\frac{1}{n_{i}}\mathbb{E}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle\leq\rho\frac{1}{\sqrt{n_{i}}}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}.\\ \end{split}$ (65)
@@
$0\geq-\rho\frac{1}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|^{2}-\frac{\lambda}{2}\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}.$ (66)
To utilize the statistical global convergence rate shown in Appendix A.3.2, we need to discuss $R$ in the same way. Recall the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ in (58):
$\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\begin{cases}\frac{C_{2}}{n}&R>\frac{1}{\sqrt{n}}\\ \frac{C_{1}}{N}+C_{2}R^{2}&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (67)
For simplicity, in the following arguments, we denote the upper bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$.
@@
$\begin{split}-\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}&=-\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\geq-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}-2\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|_{2}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}\\ &\geq-\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}-R^{2}-2R\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\end{split}$ (68)
@@
$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\mathbb{E}\left\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}+{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ \geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}-\lambda R\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}.\\ \end{split}$ (69)
Putting (67) into (69) and reorganizing, we obtain
@@
$\begin{split}0\geq&-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}+\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\mathbb{E}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}U_{0}^{2}-\frac{\lambda}{2}R^{2}-\lambda R\;U_{0}\\ \geq&\frac{\mu}{2}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\rho}{\sqrt{n_{i}}}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|_{2}-\frac{\lambda}{2}\left(U_{0}+R\right)^{2}.\end{split}$ (70)
@@
In addition, recall that in this case we set $\lambda^{\star}\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}$ in the statistical convergence analysis of the global model. If we assume data balanced, i.e. $p_{i}=\frac{1}{m}$ and $n_{i}=n,\forall i\in[m]$, it yields $\lambda^{\star}\leq\frac{\rho^{2}}{\mu nR^{2}}$ and